Write an explicit formula for the sequence -3, 1, 5, 9,...
A) an=-3+4(n-1)
B) an=-3+4n
C) an=-3+(n-1)
D) an=-3+4(n+1)
The "n" of "an" is meant to be in subscript but nothing else is.

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Write an explicit formula for the sequence -3, 1, 5, 9,...
A) an=-3+4(n-1)
B) an=-3+4n
C) an=-3+(n-1)
D) an=-3+4(n+1)
----
Procedure::
a(1) = -3
a(2) = -3+4 = 1
a(3) = -3+2*4 = 5
a(4) = -3+3*4 = 9
---
a(n) = -3+ (n-1)4
-----
Cheers,
Stan H.
--------------

Answer by solver91311(24713)



A plain text indicator for a subscript is an underscore character. a_n means . In the event you have more than one character in your subscript, enclose the subscript characters in curly braces. a_{15} means

Your question has two possible answers depending on the lower index. That is, is your sequence or ?

In the first case, answer B would give , then , and so on.

In the second case, answer A would give , then , and so on.

Not all sequences start at 1.

John

My calculator said it, I believe it, that settles it